y= sin t/ t

the task is to derive an equation for dy/dt using the quotient rule.


1
Expert's answer
2021-05-06T16:17:37-0400

Let us derive an equation for "\frac{dy}{dt}" using the quotient rule:


"\frac{dy}{dt}=\frac{t\cos t-\sin t}{t^2}"